# Reffly: Melody-Constrained Lyrics Editing Model

### v1

@@ 5 Results
5.1 Effectiveness of the heuristic
Our heuristic for identifying prominent notes and words is validated against baselines using a musician-annotated dataset of 100 song clips (§$\S$ further details in 4.1.2). The first baseline relies solely on note duration to determine prominence, similar to the decoding constraints used in Lyra Tian et al. (2023), and pairs this with our word extraction heuristic. The second baseline utilizes our note extraction heuristic but restricts prominent words to nouns and verbs.
As shown in Table 2, our prominent note extraction heuristic achieves an accuracy of 96%, substantially outperforming both baselines. Furthermore, our comprehensive heuristic for extracting prominent words and notes yields a 91% successful alignment rate  [^7] Successful alignment rate is calculated as the proportion of prominent words correctly mapped to prominent notes., surpassing the best baseline by 26%. These results underscore the effectiveness and non-trivial nature of our approach in capturing the alignment between prominent words and prominent notes (more details in Appendix E.4).
@@ 5.3 Result of Structure-Aware Generation
Among all baselines, only GPT-4 has the capability of generating full-length structured lyrics. Table 4 present the result of structure-aware generation task. Reffly demonstrated a 44% improvement on structural clarity. Figure  10 shows the generated lyrics’s clear verse-chorus-verse-chorus structure.
@@ 6 Case Study
Revision capability: Reffly rephrases sentence structures or modifies words, adding ties to ensure prominent word-note alignment. For example, in Figure 4, it rephrased ‘Your calm face reflecting vibrant colors’ into ‘eyes of peace, a canvas of hues’, and ‘You can follow my steps’ into ‘You can follow my footsteps". In both cases, the original meaning is retained. Furthermore, Lyra and SongMASS could not generate coherent lyrics. Although both Reffly and Lyra generate lyrics line-by-line, Reffly produces coherent lyrics due to our training strategy that considers the context of previously generated lines.
6.1 Ablation Study
We conduct ablation study to validate each component in Reffly. A qualitative example in Figure 7 compares different model variations. Reffly-S omits the candidate selection algorithm, and Reffly-I excludes the instruction component in training. Table 1 indicates Reffly-S perform better than Reffly-I, but is prosody score is lower comparing to Reffly. Removing both candidate selection and instruction mechanism would cause a noticeable performance drop.
@@ 7 Related work
While various M2L generators have been developed to follow control signals like music scores or themes, none provide full control over the generated content. Fan et al. (2019) employed control mechanisms to generate lyrics based on specific topics, Tian et al. (2023) used keywords and genre to control the content, and Saeed et al. (2019), used music audio to condition the generation process. Other approaches have incorporated stylistic elements, such as rhyme schemes and meter, to influence the lyrical output (Potash et al., 2015; Zhang et al., 2022c; Qian et al., 2023). Despite these efforts, the control is typically limited to broad thematic or stylistic aspects, and lacks fine-grained control such as sentence-level semantics and note-level music alignment with the score. To the best of our knowledge, Reffly is the first to provide such capabilities.
@@
LLMs have proven effective in the M2L generation task, with various attempts to integrate music representation (Zhang et al., 2022c; Lee et al., 2019; Qian et al., 2022). For example, Sheng et al. (2021) applied two transformers for cross-attention between lyrics and melody; Tian et al. (2023) considered duration of musical note in decoding stage, aligning stressed syllables with longer musical notes. However, existing models fail to align prominent words with prominent notes, resulting in poor prosody. We propose a novel heuristic for melody-lyrics alignment, achieving a 26% improvement. Reffly leverages this approach to enhance lyric generation and emotional expression.
@@ 8 Conclusion
In this work, we introduced Reffly, the first revision framework to generate high-quality song lyrics from plain text drafts while retaining their original meaning. Furthermore, to enhance the lyrics-melody alignment, we designed a heuristic to identify prominent notes in the melody and important words in lyrics. Using these features and song structure information, we developed an instruction-based mechanism to guide LLMs in generating singable lyrics with key words mapped to prominent notes. Finally, we show the diverse applicability of Reffly demonstrates diverse applicability: it generates full-fledged lyrics with specific structures from scattered user input and is the first to facilitate lyrics translation.
@@ A.2 Example of validation dataset
Figure 6: Exemplary data points in validation datasets, where experts annotate the ground truth prominent notes
@@ B.1 Qualification Task
To evaluate the Turkers’ expertise in the field, we designed a task that included the initial verse from 9 different songs, each with ground-truth labels. These songs were chosen with care to avoid unclear cases, allowing for a clear assessment of quality. The selected songs were those whose scores showed a strong correlation with the ground-truth labels. We select 49 qualified annotators out of 87 annotators, based on Pearson correlation metric. The average inter-rater agreement in therms of Pearson correlation among qualified annotators was 0.43.
@@ Appendix D Reffly Details
The candidate selection algorithm refines a list of lyric candidates to select the best match based on melody constraints, calculating ties and similarities to ensure a singable output.
Figure 9: Training data example
@@ D.2 Training Data Construction
The final piece of input data is music constraint. Due to the lack of melody-aligned data, we generate a pseudo music constraint for lyric dataset, independent of its melody. This approach addresses both the scarcity and copyright issues associated with aligned data. We assume that the lyrics in the training data exhibit good prosody and are singable. we assign a special token, symbolizing a pseudo note, to each syllable in the sentence (singability assumption). If the syllable associated with the special token occupies the stress position of a prominent word, the token denotes an prominent note, otherwise it represents a less prominent note (good-prosody assumption).
@@ E.2 Prominent note extraction heuristic details
E.3 Prominent words extraction heuristic details
We identify all non-stop words whose part-of-speech are noun, verbs, and adjectives as prominent words.
E.4 Results for heuristics

### v2

@@ 5 Results
5.1 Effectiveness of the proposed heuristic
Our heuristic for identifying prominent notes and words is validated against baselines using a musician-annotated dataset of 100 song clips (§$\S$ further details in 3.1). The first baseline (Duration-only) relies solely on note duration to determine prominence, similar to the decoding constraints used in Lyra Tian et al. (2023), and pairs this with our word extraction heuristic. The second baseline (Comprehensive w/o adj.) utilizes our note extraction heuristic but restricts prominent words to nouns and verbs.
As shown in Table 2, our prominent note extraction heuristic achieves an accuracy of 96%, substantially outperforming both baselines. Furthermore, our comprehensive heuristic for extracting prominent words and notes yields a 91% alignment success rate [^7] Alignment success rate is the accuracy of prominent words correctly mapped to prominent notes; note extraction success rate is the accuracy of extracting prominent notes, surpassing the best baseline by 26%. These results underscore the effectiveness and non-trivial nature of our approach in capturing the alignment between prominent words and prominent notes (more details in Appendix E.3).
@@ 5.2 Result of Lyrics Generation from Arbitrary Content
5.3 Ablation Study
We conduct an ablation study to validate each component in Reffly. A qualitative example in Figure 7 (Appendix) compares variations of the model. Candidate selection algorithm select optimal candidate generated from revision model (see Figure 2). Instruction component in training simplifies the task for the revision module and enabling better context awareness. As shown in Table 1, removing the candidate selection (Reffly w/o S.) or instruction mechanisms (Reffly w/o I.) results in a noticeable performance decline in almost every metrics compared with Reffly.
@@ 5.4 Result of Structure-Aware Generation
Among all baselines, only GPT-4 has the capability of generating full-length structured lyrics. Table 4 shows the results for the structure-aware generation task, where Reffly achieved a 44% improvement in structural clarity. Figure  10 in Appendix shows the generated lyrics’s clear verse-chorus-verse-chorus structure.
@@ 6 Case Study
Revision capability: Reffly rephrases sentence structures or modifies words, adding ties to ensure prominent word-note alignment. For example, in Figure 4, it rephrased ‘Your calm face reflecting vibrant colors’ into ‘eyes of peace, a canvas of hues’, and ‘You can follow my steps’ into ‘You can follow my footsteps’. In both cases, the original meaning is retained. Although both Reffly and Lyra generate lyrics line-by-line, Reffly produces coherent lyrics due to our training strategy that considers the context of previously generated lines.
@@ 7 Related work
While various lyric generators and AI-assisted lyrics writing system have been developed to follow control signals like themes, rhyme, specific text format, or in-filling text template (Ma et al., 2020; Li et al., 2020; Liu et al., 2020; Zhang et al., 2022c; Sun et al., 2022), none provide full control over the sentence-level semantics of generated content and follow melody constraint; Fan et al. (2019) employed control mechanisms to generate lyrics based on specific topics, Tian et al. (2023) used keywords and genre to control the content, and Saeed et al. (2019), used music audio to condition the generation process. Other approaches have incorporated stylistic elements, such as rhyme schemes and meter or text format to influence the lyrical output (Potash et al., 2015; Zhang et al., 2022c). Despite these efforts, they have insufficient sentence-level semantics and note-level music alignment. Reffly can generate coherent, full-fledged high quality melody-aligned lyrics with sentence-level semantics control.
@@
LLMs have proven effective in the M2L generation task, with various attempts to integrate music representation (Lee et al., 2019; Qian et al., 2022). For example, Sheng et al. (2021) applied two transformers for cross-attention between lyrics and melody; Tian et al. (2023); Qian et al. (2023) considered duration of musical note and stressed syllables mapping, beat, or song structures during lyrics generation. ReLyMe (Zhang et al., 2022a), a lyrics-to-melody (L2M) generation model, considers the mapping of keywords to stressed beat locations but does not discuss how pitch influences the determination of prominent notes or how keywords are identified. While revision approaches have been explored in poetry generation (Andrea et al., 2021), REFFLY is the first framework for melody-constrained lyrics revision with sentence-level semantic control. Furthermore, existing M2L models fail to align prominent words with prominent notes, resulting in poor prosody. We propose a novel heuristic for melody-lyrics alignment, achieving a 26% improvement. Reffly enhances lyric generation and emotional expression.
@@ 8 Conclusion
We introduced Reffly, the first melody constrained revision framework to generate high-quality lyrics from plain text drafts while retaining original meaning. To enhance the lyrics-melody alignment, we designed a heuristic to identify and align prominent notes and words. Finally, we show Reffly excel across diverse applicability.
@@ A.2 Example of validation dataset
Figure 6: Exemplary data points in validation datasets, where experts annotate the ground truth prominent notes, actual data points have at three to five musical phrases.
@@ B.1 Qualification Task
To evaluate the Turkers’ expertise in the field, we designed a task that included the initial verse from 9 different songs, each with ground-truth labels. These songs were chosen with care to avoid unclear cases, allowing for a clear assessment of quality. The selected songs were those whose scores showed a strong correlation with the ground-truth labels. We select 49 qualified annotators out of 87 annotators, based on Pearson correlation metric. The average inter-rater agreement interms of Pearson correlation among qualified annotators in qualification task was 0.43.
@@ C.3 Computation cost
C.4 More details regarding to experiment metrics
We utilized the GPT-2 Large model, which has 774 million parameters, to calculate perplexity. Prominent word-note matching rate, or matching rate in Table 1, is the accuracy of stressed syllables of prominent words being correctly mapped to prominent notes.
@@ Appendix D Reffly Details
The candidate selection algorithm refines a list of lyric candidates to select the best match based on melody constraints, calculating ties and similarities to ensure a singable output. Note that there is no need for human to post process the lyrics.
Figure 9: Training data example.
@@ D.2 Training Data Construction
The final piece of input data is music constraint. Due to the lack of melody-aligned data, we generate a pseudo music constraint for lyric dataset, independent of its melody. This approach addresses both the scarcity and copyright issues associated with aligned data. We assume that the lyrics in the training data exhibit good prosody and are singable. we assign a special token, symbolizing a pseudo note, to each syllable in the sentence (singability assumption). If the syllable associated with the special token occupies the stress position of a prominent word, the token denotes an prominent note, otherwise it represents a less prominent note (good-prosody assumption). This way, we “back-translating" (Longshen et al. (2023)) the pseudo music constraint from pure lyric-side data.
@@
We used the default LoRA implementation from the official LLaMA2-13b GitHub repository to fine-tune the model. The revision module was trained for 3 epochs using a dataset of 3,500 data points.
@@ E.2 Prominent note extraction heuristic details
We also provide the mathmatical formulations for prominent note extraction:
1. Time signature:
As shown in table 5, a function can determine if a note is in an important location in terms of time signature. Suppose the time signature for this melody is $T$, and the corresponding list for stressed location is $SL_{T}$
$\text{stressed}(n_{i})=\begin{cases}1&\text{if }\text{if offset}(n_{i})\text{ in }SL_{T}\\ 0&\text{, otherwise}\end{cases}$ (2)
2. Rhythmic type:
We implemented two simple rules:
1) for a given k consecutive notes $n_{i},...n_{i+k}$ that are connected by one tie, we replace $n_{i},...n_{i+k}$ as a new note $n_{i}^{\prime}$, and the duration for $n_{i}^{\prime}$ is
$duration(n_{i}^{\prime})=\sum_{j=0}^{k}duration(m_{i+j})$,
and offset (beginning of $n_{i}^{\prime}$) is:
$offset(n_{i}^{\prime})=offset(n_{i})$
2) After combining all notes that connected by tie, we check syncopation. If a note is in a weak location but its duration is longer than previous note, it is a syncopation.
3. Pitch:
A note is more important if there as a dramatic change in pitch compared to previous note. Based on this assumption, we have
$\text{jump}(n_{i})=\begin{cases}1&\text{if }\Delta\text{pitch}(n_{i})>AIJ,\\ 0&\text{otherwise},\end{cases}$ (3)
where
$\Delta\text{pitch}(n_{i})$ $=|\text{pitch}(n_{i})-\text{pitch}(n_{i-1})|,$
$AIJ$ $=\frac{\sum_{i=1}^{x}\Delta\text{pitch}(n_{i})}{x},$
and $\text{and }x$ is the number of notes.
Collectively, the importance of note $m_{i}$ is defined by
$\text{M}(m_{i})=\begin{cases}1&\text{if }\text{if stressed, jump, or syncopate}(m_{i})\\ 0&\text{, otherwise}\end{cases}$ (4)
, where (0)1 means the note is an (un)important note.
E.3 Results for heuristics
@@
Note that in the heuristic, we are consider the stressed syllable position by aligning stressed syllables of prominent words to prominent notes.
@@
E.4 An illustrative example of how the tie is added
In figure 4, we need to add two ties to the first sentence “eyes of peace a canvas of hues”, because there are two more notes in the melody than number of syllables. Here, we discuss why there is a tie added to “Do” instead of “Mi”. This is because at here, “Mi” would be identified as a prominent note (because it is in a stressed beat in 4/4) and “peace” as a prominent word (because it is a noun). Note that here “Do” is not a prominent note, because it is neither in a stressed location, nor a syncopated note, nor a note with big pitch jump from the previous note. To maximize the number of prominent words mapped to prominent notes, a tie would be added at “Do”, and the important word “peace” is corresponding to the prominent note “Mi”. This process is detailed at Step 3.2 (§$\S$3.3) and Algorithm 1, candidate selection algorithm. This entire process is handled algorithmically, without requiring human inspection.